# 1Max-Planck-Institut für Extraterrestrische Physik, Postfach 1312, Giessenbachstr., D-85748 Garching, Germany 2Aix Marseille Univ, Université de Toulon, CNRS, CPT, Marseille, France 3Instituto de Astronomía Teórica y Experimental (CONICET-UNC), Laprida 854, X5000BGR, Córdoba, Argentina

### v1

@@ Abstract
We extend the evolution mapping approach, originally proposed by Sánchez et al. (2022) to describe non-linear matter density fluctuations, to statistics of the cosmic velocity field. This framework classifies cosmological parameters into shape parameters, which determine the shape of the linear matter power spectrum, $P_{\rm L}(k,z)$, and evolution parameters, which control its amplitude at any given redshift. Evolution mapping leverages the fact that density fluctuations in cosmologies with identical shape parameters but different evolution parameters exhibit remarkably similar non-linear evolutions when expressed as a function of the clustering amplitude. We use a suite of N-body simulations sharing identical shape parameters but spanning a wide range of evolution parameters. Using an efficient method for estimating the volume-weighted velocity field based on the Voronoi tesselation of the simulation particles, we study the non-linear evolution of the power spectra of the velocity divergence, $P_{\theta\theta}(k)$, and its cross-power spectrum with the density field, $P_{\delta\theta}(k)$. By analysing snapshots at redshifts where the linear matter perturbations have the same amplitude, we demonstrate that evolution mapping accurately applies to $P_{\theta\theta}(k)$ and $P_{\delta\theta}(k)$. Deviations at small scales can be modelled in terms of differences in the suppression factor, $g(a)=D(a)/a$, akin to those observed for the density field. Evolution mapping simplifies the description of the cosmological dependence of non-linear density and velocity statistics, streamlining the sampling of large cosmological parameter spaces for the analysis of cosmological observables.
@@
Cosmology: large-scale structure of Universe – Methods: numerical
@@ 1 Introduction
N-body simulations are one of the most precise tools we use nowadays for detailed studies of non-linear structure formation (Efstathiou et al., 1985; Kravtsov et al., 1997; Teyssier, 2002; Springel, 2005; Springel et al., 2021; Potter et al., 2017; Garrison et al., 2019). By numerically solving the equations of motion for a large number of particles, these simulations can model the evolution of cosmic structures from the early universe to the present day. N-body simulations are crucial for testing theoretical models against observational data, as they can account for complex interactions and nonlinear gravitational effects that are difficult to capture analytically. However, such accurate simulations come at the cost of high computational complexity. Large suites of simulations are often needed to explore different cosmologies, requiring numerous CPU hours on high-performance computer clusters. For this reason, exploiting our understanding of cosmology is paramount to aiding the exploration of wide cosmological parameter spaces.
With this in mind, Sánchez et al. (2022) pointed out that a useful degeneracy is revealed if one classifies cosmological parameters into two groups: shape parameters, $\Theta_{\rm s}$, which determine the shape of the linear matter power spectrum, $P_{\rm L}(k,z)$, and evolution parameters, $\Theta_{\rm e}$, which only affect its amplitude at any given redshift. With this definition, the time evolution of $P_{\rm L}(k,z)$ in models sharing the same shape parameters but different evolution parameters can be mapped from one to the other by relabelling the redshifts that correspond to the same clustering amplitude. This evolution mapping relation not only simplifies comparing different cosmological models but also accurately describes the non-linear matter power spectrum or the full density field (Sánchez et al., 2022).
However, the density field cannot be observed directly. In particular, galaxy positions along the line of sight can only be inferred from their redshift, which is also influenced by the galaxies’ peculiar velocities (Kaiser, 1987). The modelling of this effect—the so-called redshift-space distortions (RSD)—is fundamental for producing accurate predictions of the observational data. Moreover, these velocities arise from gravitational interactions within the large-scale structure, providing additional information about the underlying mass distribution and the dynamics of cosmic structures (Guzzo et al., 2008). For this reason, this paper aims to use evolution mapping to ease the modelling of velocity statistics.
@@
We begin in section 2 by summarising the results of Sánchez et al. (2022) on the matter power spectrum and argue why these should also apply to $P_{\theta\theta}(k)$ and $P_{\delta\theta}(k)$. In section 3, we present the Aletheia simulations, and in section 4, we describe the method used to estimate the velocity field from them. In section 5 we present and discuss our results. Finally, we give our conclusions in section 6. We include in appendix A a discussion on a problem that might emerge when dealing with Gadget-4 simulations if the user is interested in the velocities of snapshots at a very precise redshift.
@@ 2.1 The matter power spectrum
One of the successes of the standard $\Lambda$CDM model resides in its ability to describe the physics of cosmic structure formation through a relatively small number of parameters. Although different parameter bases are equivalent, choosing one that explicitly exhibits degeneracies in the statistics of interest (in our case, the matter power spectrum) can ease its modelling and understanding. With this in mind, we focus on the parameters
@@
As evolution parameters only affect one degree of freedom in $P_{\rm L}(k,z)$ —its amplitude— they are perfectly degenerate among each other and with the redshift $z$. The impact of all such parameters can thus be summarized in a single quantity that captures the amplitude of $P_{\rm L}(k,z)$. This is often given in terms of the RMS variance of the linear-theory density field at a reference scale $R$ as
@@
The perfect mapping enforced by eq. 7 holds only at the linear level since, in the non-linear dynamics, we cannot treat the evolution of each $k$ mode separately anymore. However, since $P_{\rm L}(k,z)$ is the main driver of the non-linear matter power spectrum, the degeneracy exposed in eq. 7 is inherited to a good extent also in the non-linear regime. In particular, Sánchez et al. (2022) showed that in N-body simulations of cosmologies with the same shape parameters but different evolution parameters, if one evaluates the matter power spectra at redshifts at which the different cosmologies have the same $\sigma_{12}$, such power spectra are extremely similar. This provides an approximate evolution mapping relation
@@ 2.2 The cosmic velocity field
where $t$ is the cosmic time associated to a comoving observer, x are the comoving coordinates and $\textbf{v}\coloneqq a\mathrm{d}\textbf{x}/\mathrm{d}t$ is the peculiar velocity relative to the background expansion. Turning into Fourier space, the relation of eq. 11 can be expressed as
@@ 3 The Aletheia simulations
We produced five snapshots for each model at the redshifts at which each cosmology reaches a given value of $\sigma_{12}$. This means that our snapshots are at the same $\sigma_{12}$ rather than at the same redshift, and thus, all these snapshots have the same $P_{\rm L}(k,z)$ (and consequently also identical linear-theory $P_{\theta\theta}(k,z),P_{\delta\theta}(k,z)$, eq. 15). To obtain the snapshots at the exact required redshift, we adopt a start-and-stop approach for running our simulations with Gadget-4, ensuring that our snapshots only contain fully synchronized particles at the desired redshift. Even though this solution disrupts the organised time domain decomposition in Gadget-4, we found that it gives sufficiently accurate results for this work. Alternatively, one could adopt the option OUTPUT_NONSYNCHRONIZED_ALLOWED, which provides snapshots with synchronized positions but unsynchronized velocities. Although one can mitigate the effect of unsynchronized velocities by reducing the maximum allowed time step, we advocate against using this option for its impact on velocity statistics, which is often relevant and difficult to estimate a priori. We show in appendix A different tests that highlight the impact of the usage of this flag on velocity statistics.
The simulations in this work replicate the ones employed in Sánchez et al. (2022), with some differences. The most important one is the adoption of the start-and-stop method in place of unsynchronized velocities. Because of the inaccuracies induced by the latter method in the velocity power spectra, we decided to perform again the whole suite of simulations, avoiding using the aforementioned option. Note that this did not impact the particle positions; hence, the $P_{\delta\delta}(k)$ in our simulations are the same as the ones in the old set. Secondly, the simulations of Sánchez et al. (2022) were performed starting from fixed redshift rather than fixed $\sigma_{12}$ as described above. Finally, we replaced the Early Dark Energy (EdE) model with the EdS one. This choice comes from the fact that the EdE simulation was run with an effective model that only considered the impact of EdE on the Hubble function and the growth factor. More generally, EdE models can also change the shape of the transfer function. Thus, we decided to substitute that model with an EdS cosmology, which is often used as a baseline for theoretical modelling and will be useful for future applications.
@@ 4 Estimating the density and velocity power spectra
In this work, we follow a different approach based on the work of Bernardeau & van de Weygaert (1996), that implies a lower-order approximation of v. We assume that the velocity field is constant inside each Voronoi cell [^2] In a Voronoi tesselation, space is divided into regions around a set of particles such that each region contains all the points closer to its corresponding particle than to any other, forming a set of convex polyhedra. of the particles in the simulation. With this piece-wise approximation, the field is discontinuous at the edges of each cell and some small-scale information is lost. However, we propose a Monte-Carlo variation of this Voronoi method that compensates a small loss in accuracy by considerably reducing the number of calculations needed. This approach can be implemented without calculating the full Voronoi tesselation of our simulations. We simply populate the boxes with a uniform set of points, and assign to each of them the velocity of the closest N-body particle. This results in a new particle distribution that reproduces our piece-wise approximation of the velocity field. The now uniformly distributed samples can then be used to estimate the smoothed velocity field by means of commonly used mass assignment schemes. Hereafter, we refer to this approach as the Monte-Carlo Voronoi method (MC-Voronoi).
We start by populating the simulation box with a sample of particles following a glass-like distribution generated using the recipe of Dávila-Kurbán et al. (2021). These samples provide a more homogeneous covering of the volume than a Poisson distribution with the same number of points, reducing the shot-noise contribution in our estimates. Each point in this distribution is assigned a velocity matching that of its closest neighbour among the simulation particles. We then use this new set of points to compute the CIC smoothed velocity field, $\bar{\bf v}$, on a regular mesh
@@ 5 Results
In this section, we present the power spectra estimated with the technique described in section 4. fig. 1 shows the $P_{\theta\theta}(k)$ and $P_{\delta\theta}(k)$ estimated from the Aletheia simulations in the left and right panels, respectively. The solid lines correspond to the results for the cosmologies described in table 2, while the dashed lines indicate the linear theory predictions. As expected, given that the snapshots of the different cosmologies correspond to the same values of $\sigma_{12}$, all models have the same power spectra at large scales but present some small deviations at small scales where the evolution mapping relation becomes approximate. The measurements of $P_{\delta\theta}(k)$ show a steep drop and a change of sign (which is not visible in the figure) that correspond to the appearance of vorticity. While the velocity field has no vorticity at the linear level, it is produced by the non-linear evolution of the density fluctuations (Pueblas & Scoccimarro, 2009; Hahn et al., 2015). Such vorticity disrupts the cosmic flow and breaks the tight correlation between the density and velocity fields. Hence, vorticity causes the cross-power spectrum to drop at the scales when it becomes important and then change its sign at the scale when shell crossing happens.
@@ 6 Conclusions
We advocate against the usage of non-synchronized outputs in Gadget-4. Even though this option is not the default one, users may be tempted to activate it when needing snapshots at a specific redhsift. Although this results in correct particle positions, velocities can be biased in ways that are difficult to predict. We present a series of illustrative tests of this effect in appendix A. We thus suggest to rather use a start-stop technique from snapshot to snapshot to ensure that each of them contains fully synchronised particles at the desired redshift.
@@ Acknowledgements
We would like to thank Raul Angulo and the BACCO project for providing useful scripts for benchmarking the codes used in this work. We also would like to thank Sofia Contarini, Carlos Correa, Andrea Fiorilli, Luca Fiorino, Nelson Padilla, Alejandro Perez, Andrea Pezzotta, Agne Semenaite and Matteo Zennaro for their help and useful discussions. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC 2094 – 390783311. The Aletheia simulations were carried out and post-processed on the HPC system Cobra of the Max Planck Computing and Data Facility (MPCDF) in Garching, Germany.

### v2

@@ Abstract
We extend the evolution mapping approach, introduced in the first paper of this series to describe non-linear matter density fluctuations, to statistics of the cosmic velocity field. This framework classifies cosmological parameters into shape parameters, which determine the shape of the linear matter power spectrum, $P_{\rm L}(k,z)$, and evolution parameters, which control its amplitude at any redshift. Evolution mapping leverages the fact that density fluctuations in cosmologies with identical shape parameters but different evolution parameters exhibit similar non-linear evolutions when expressed as a function of clustering amplitude. We analyse a suite of N-body simulations sharing identical shape parameters but spanning a wide range of evolution parameters. Using a method for estimating the volume-weighted velocity field based on the Voronoi tesselation of simulation particles, we study the non-linear evolution of the velocity divergence power spectrum, $P_{\theta\theta}(k)$, and its cross-power spectrum with the density field, $P_{\delta\theta}(k)$. We demonstrate that the evolution mapping relation applies accurately to $P_{\theta\theta}(k)$ and $P_{\delta\theta}(k)$. While this breaks down in the strongly non-linear regime, deviations can be modelled in terms of differences in the suppression factor, $g(a)=D(a)/a$, similar to those for the density field. Such modelling describes the differences in $P_{\theta\theta}(k)$ between models with the same linear clustering amplitude to better than 1 percent accuracy at all scales and redshifts considered. Evolution mapping simplifies the description of the cosmological dependence of non-linear density and velocity statistics, streamlining the sampling of large cosmological parameter spaces for cosmological analysis.
@@
cosmology: cosmological parameters — cosmology: large-scale structure of the Universe — cosmology: theory — methods: numerical — software: simulations
@@ 1 Introduction
N-body simulations are the ideal tool for detailed studies of non-linear structure formation (Efstathiou et al., 1985; Kravtsov et al., 1997; Teyssier, 2002; Springel, 2005; Springel et al., 2021; Potter et al., 2017; Garrison et al., 2019). By numerically solving the equations of motion for a large number of particles, these simulations can model the evolution of cosmic structures from the early universe to the present day. N-body simulations are crucial for testing theoretical models against observational data, as they can account for nonlinear gravitational effects that are difficult to capture analytically. However, such accurate simulations come at the cost of high computational complexity. Large suites of simulations are often needed to explore different cosmologies, requiring numerous CPU hours on high-performance computer clusters. For this reason, exploiting our understanding of cosmology is paramount to aiding the exploration of wide cosmological parameter spaces.
With this in mind, Sánchez et al. (2022) pointed out that a useful degeneracy is revealed if one classifies cosmological parameters into two groups: shape parameters, $\Theta_{\rm s}$, which determine the shape of the linear matter power spectrum, $P_{\rm L}(k,z)$, and evolution parameters, $\Theta_{\rm e}$, which only affect its amplitude at any given redshift. With this definition, the time evolution of $P_{\rm L}(k,z)$ in models sharing the same shape parameters but different evolution parameters can be mapped from one to the other by relabelling the redshifts that correspond to the same linear clustering amplitude. This evolution mapping relation not only simplifies comparing different cosmological models but also accurately describes the non-linear matter power spectrum of the full density field (Sánchez et al., 2022).
However, the density field cannot be observed directly. In particular, galaxy positions along the line of sight can only be inferred from their redshift, which is also influenced by the galaxies’ peculiar velocities (Kaiser, 1987). The modelling of this effect—the so-called redshift-space distortions (RSD)—is fundamental for producing accurate predictions of the observational data. Moreover, these velocities arise from gravitational interactions within the large-scale structure, providing additional information about the underlying mass distribution and the dynamics of cosmic structures (Guzzo et al., 2008). For this reason, this paper aims to use evolution mapping to simplify the modelling of velocity statistics.
@@
We begin in section 2 by summarising the results of Sánchez et al. (2022) on the matter power spectrum and argue why these should also apply to $P_{\theta\theta}(k)$ and $P_{\delta\theta}(k)$. In section 3, we present the Aletheia simulations, and in section 4, we describe the method used to estimate the velocity field from them. In section 5 we present and discuss our results. Finally, we give our conclusions in section 6. We include in appendix A a discussion on a problem that emerges when dealing with Gadget-4 simulations if the user is interested in the velocities of snapshots at a very precise redshift.
@@ 2.1 The matter power spectrum
One of the successes of the standard $\Lambda$CDM model resides in its ability to describe the physics of cosmic structure formation through a relatively small number of parameters. Although different parameter bases are equivalent, choosing one that explicitly exhibits degeneracies in the statistics of interest (in our case, the matter power spectrum) can simplify its modelling and understanding. With this in mind, we focus on the parameters
@@
Since changes in the values of the evolution parameters, as well as in the redshift $z$ (through the growth factor), only affect the amplitude of $P_{\rm L}(k,z)$, they are all degenerate with each other. Therefore, the effects of these parameters can be captured by a single quantity representing the amplitude of $P_{\rm L}(k,z)$. This is often given in terms of the RMS variance of the linear-theory density field at a reference scale $R$ as
@@
The perfect correspondence described by eq. 7 is strictly valid only in the linear regime, as non-linear dynamics couples the evolution of different Fourier modes. Nevertheless, since $P_{\rm L}(k,z)$ plays a dominant role in shaping the non-linear matter power spectrum, the degeneracy outlined in eq. 7 largely carries over to the non-linear regime. This effect is especially noticeable in the framework of standard perturbation theory (SPT). If we assume that the perturbation kernels are independent of cosmological parameters —which holds for an Einstein-de-Sitter (EdS) universe and is a reasonable approximation even for non-standard cosmologies (Takahashi, 2008; Taruya, 2016; Garny & Taule, 2021)— SPT predicts that the non-linear power spectrum depends solely on $P_{\rm L}(k,z)$, without explicit dependence on the cosmological parameters. Sánchez et al. (2022) used N-body simulations of cosmologies sharing the same shape parameters but differing in evolution parameters to show that the matter power spectra measured at redshifts where these cosmologies have the same value of $\sigma_{12}$ remain extremely similar. This provides an approximate evolution mapping relation
@@ 2.2 The cosmic velocity field
where $t$ is the cosmic time associated to a comoving observer, x are the comoving coordinates and $\textbf{v}\coloneqq a\mathrm{d}\textbf{x}/\mathrm{d}t$ is the peculiar velocity relative to the background expansion. Going to Fourier space, eq. 11 can be expressed as
@@ 3 The Aletheia simulations
We produced five snapshots for each model at the redshifts at which each cosmology reaches a given value of $\sigma_{12}$. This means that our snapshots are at the same $\sigma_{12}$ rather than at the same redshift, and thus, all these snapshots have the same $P_{\rm L}(k,z)$ (and consequently also identical linear-theory $P_{\theta\theta}(k,z),P_{\delta\theta}(k,z)$, eq. 15).
The simulations in this work replicate the ones employed in Sánchez et al. (2022), with some differences. First, we replaced the Early Dark Energy (EdE) model with the EdS one. This choice comes from the fact that the EdE simulation was performed with an effective model that only considered the impact of EdE on the Hubble function and the growth factor. More generally, EdE models can also change the shape of the transfer function. Thus, we decided to substitute that model with an EdS cosmology, which is often used as a baseline for theoretical modelling and will be useful for future applications. Secondly, the simulations of Sánchez et al. (2022) were performed starting from fixed redshift rather than fixed $\sigma_{12}$ as described above.
Finally, the most important change resides in the way the Gadget-4 snapshots are produced. In particular, the simulations in (Sánchez et al., 2022) were performed with a compile-time option (OUTPUT_NONSYNCHRONIZED_ALLOWED) to address a limitation in how Gadget-4 produces snapshots. As Gadget-4 uses an adaptive time-stepping scheme, it synchronizes particles only at specific times, known as synchronization points. This means particles in different regions of the simulation (dense or underdense) are updated at different rates based on their local conditions. As a result, particles are only synchronized at these designated points. When the user requests a snapshot at a given redshift, Gadget-4’s default behaviour is to save it at the closest available synchronization point. This may result in small shifts in redshift, undesirable in our analysis in which we want to compare simulations at the exact same value of $\sigma_{12}$. The OUTPUT_NONSYNCHRONIZED_ALLOWED option allows Gadget-4 to output snapshots at the exact redshift requested by linearly drifting particle positions to the desired time. However, while the positions are adjusted to the correct redshift, the particle velocities are left unchanged, meaning they correspond to different times. As a result, the snapshot contains inconsistent velocities at different redshifts, which can introduce errors in velocity-related analyses.
To address this problem while ensuring that the snapshots are exactly at the desired redshifts, we performed this new set of simulations with a start-stop method. This method leverages the fact that all particles are naturally synchronized at the start and end of each simulation run. By stopping the simulation at each required snapshot and restarting it with adjusted start and end times, we ensure that all particles are fully synchronized at the exact redshift requested. This guarantees consistent outputs for both positions and velocities. Readers interested in a more detailed discussion of the problem can find it in appendix A, along with tests of the start-stop method.
@@ 4 Estimating the density and velocity power spectra
In this work, we follow a different approach based on the work of Bernardeau & van de Weygaert (1996), that implies a lower-order approximation of v. We assume that the velocity field is constant inside each Voronoi cell [^2] In a Voronoi tesselation, space is divided into regions around a set of particles such that each region contains all the points closer to its corresponding particle than to any other, forming a set of convex polyhedra. of the particles in the simulation. With this piece-wise approximation, the field is discontinuous at the edges of each cell and some small-scale information is lost. However, we propose a Monte-Carlo variation of this Voronoi method that compensates a small loss in accuracy by considerably reducing the number of calculations needed. This approach can be implemented without calculating the full Voronoi tesselation of our simulations. We simply populate the boxes with points sampled from a distribution that evenly covers the volume (in our case, a glass-like distribution, see below), and assign to each of them the velocity of the closest N-body particle. This results in a new particle distribution that reproduces our piece-wise approximation of the velocity field. The now evenly distributed samples can be used to estimate the smoothed velocity field by means of commonly used mass assignment schemes. Hereafter, we refer to this approach as the Monte-Carlo Voronoi method (MC-Voronoi).
In our implementation, we start by populating the simulation box with a sample of particles following a glass-like distribution generated using the recipe of Dávila-Kurbán et al. (2021). These samples provide a more homogeneous covering of the volume than a Poisson distribution with the same number of points, reducing the shot-noise contribution in our estimates. Each point in this distribution is assigned a velocity matching that of its closest neighbour among the simulation particles. We then use this new set of points to compute the CIC smoothed velocity field, $\bar{\bf v}$, on a regular mesh
@@ 5 Results
In this section, we present the power spectra estimated with the technique described in section 4. fig. 1 shows the $P_{\theta\theta}(k)$ and $P_{\delta\theta}(k)$ estimated from the Aletheia simulations in the left and right panels, respectively. The solid lines correspond to the results for the cosmologies described in table 2, while the dashed lines indicate the linear theory predictions. As expected, given that the snapshots of the different cosmologies correspond to the same values of $\sigma_{12}$, all models have the same power spectra at large scales but present some small deviations at small scales where the evolution mapping relation becomes approximate. The measurements of $P_{\delta\theta}(k)$ show a steep drop and a change of sign (which is not visible in the figure) that correspond to the appearance of vorticity. While the velocity field has no vorticity at the linear level, it is produced by the non-linear evolution of the density fluctuations (Pueblas & Scoccimarro, 2009; Hahn et al., 2015). Such vorticity disrupts the cosmic flow and breaks the tight correlation between the density and velocity fields. Hence, vorticity causes the cross-power spectrum to drop at the scales when it becomes important and then change its sign at the scale of shell crossing.
@@ 6 Conclusions
We highlight a limitation in Gadget-4 when performing velocity-related analyses with simulation snapshots that require a specific redshift. The OUTPUT_NONSYNCHRONIZED_ALLOWED option, while providing accurate particle positions at exact redshifts, leads to inconsistencies in velocities because particles are not synchronized in time, introducing potential biases that are difficult to predict. We thus recommend using a start-stop method instead, which we discuss in more detail in appendix A where we address the problem.
@@ Acknowledgements
We would like to thank Raul Angulo and the BACCO project for providing useful scripts for benchmarking the codes used in this work. We also would like to thank Sofia Contarini, Carlos Correa, Andrea Fiorilli, Luca Fiorino, Nelson Padilla, Alejandro Perez, Andrea Pezzotta, Agne Semenaite and Matteo Zennaro for their help and useful discussions. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC 2094 – 390783311. The Aletheia simulations were carried out and post-processed on the HPC system Cobra of the Max Planck Computing and Data Facility (MPCDF) in Garching, Germany. We acknowledge support from the European Research Executive Agency HORIZON-MSCA-2021-SE-01 Research and Innovation programme under the Marie Sklodowska-Curie grant agreement number 101086388 (LACEGAL).